D. 3 m/s 10. A spring gun is used to project a 106 g mass horizontally across a frictionless surface. The spring is compressed 4.5 cm and has a spring constant of 33 N/m. What is the initial magnitude of the acceleration of the mass the nstant the spring gun fires? A. 140 m/s2 B. 14 m/s2 C. 28 m/s2 D. 1.4 m/s2 11. A kid of mass 20 kg is at the center of a playground merry-go-round that is rotating without friction at 18 rev/min. The mass of the merry-go-round is 160 kg, its diameter is 3.50 m, and its moment of inertia is Ve mr2. The kid can be treated as a point mass (l mr2). The kid slowly crawls to the rim of the merry-go-round. Find the new angular speed in rev/min. A. 17.8 rev/min B. 14.4 rev/min C. 105 rev/min D. 1.50 rev/min
Answer #1

10.

Considering hooke's law

ma = kx

a = kx/m = 33* 4.5*10^-2 / 0.106

a = 14 m/s^2

======

11.

Considering conservation of angular momentum

(0.5 M r^2 ) w1 = (0.5 Mr^2 + mr^2) w2

0.5*160* 3.5^2* 18 = (0.5* 160*3.5^2 + 20* 3.5^2) w2

w2 = 14.4 rev/min

=====

Do comment in case any doubt will reply for sure . Goodluck